OpenStudy (anonymous):

what conversion factors should be used to convert 18 mi/hr to ft/sec? what conversion factors should be used to convert 18 mi/hr to ft/sec? @Mathematics

OpenStudy (asnaseer):

1 mile = 5280 feet 1 hour = 60 seconds so if you travel 18 miles in one hour - that means you travelled (18*5280) feet in one hour - which means you travelled (18*5280)/60 feet in one second so conversion factor is 5280/60 = 88

OpenStudy (asnaseer):

oops! 1 hour = 3600 seconds

OpenStudy (asnaseer):

please correct my mistake when you do the calculation

OpenStudy (anonymous):

so 5280/3600?

OpenStudy (asnaseer):

yes, which works out at 22/15

OpenStudy (anonymous):

okay thank you! & do you know how to determine if each rate is a unit rate?

OpenStudy (asnaseer):

I am not sure I understand that question

OpenStudy (anonymous):

like..

OpenStudy (anonymous):

3 liters per 60 seconds

OpenStudy (asnaseer):

ah - ok - just looked up what that means, see: http://www.aaamath.com/rat-unit-rate.htm so yes, this is a unit rate just like the original miles per hour was as well

OpenStudy (anonymous):

the problem is, i have no idea how to solve these. :/

OpenStudy (asnaseer):

by the way - 3 liters per 60 seconds would NOT be a unit rate but I think 3 liters per minute would be unless I have misunderstood the term

OpenStudy (asnaseer):

if you break the problem down into simple steps (as I did above) then you should be able to grasp it

OpenStudy (anonymous):

that's what it says on the problem.

OpenStudy (anonymous):

okayy.:)

OpenStudy (anonymous):

but can you do this one so like i can have an example? :)

OpenStudy (asnaseer):

so 18 miles per hour = (18*22/15) feet per second = 26.4 ft/sec